Determine the force of gravitational attraction between the earth (m = 5.98 x 1024 kg ) and a 70-kg physics student if the student is in an airplane at 40000 feet above earth’s surface. This would place the student a distance of 6.39 x 106 m from earth’s center.


1
Expert's answer
2021-03-17T16:52:56-0400
"F=\dfrac{GMm_1}{R^2},""F=\dfrac{6.67\cdot10^{-11}\ \dfrac{Nm^2}{kg^2}\cdot5.98\cdot10^{24}\ kg\cdot70\ kg}{(6.39\cdot10^6\ m)^2}=684\ N."